OpenStudy (anonymous):

A 10-ft diagonal brace on a bridge connects a support of the center of the bridge to a side support on the bridge. The horizontal distance that it spans is 2 ft longer than the height that it reaches on the side of the bridge. Find the horizontal and vertical distances spanned by this brace.

OpenStudy (anonymous):

Not enough info...?

OpenStudy (anonymous):

thats all i have

OpenStudy (anonymous):

we have been working on factoring if that helps

OpenStudy (anonymous):

polynomials

OpenStudy (anonymous):

Well, I have an image of your bridge. The brace extends 2 feet past where it passes the bridge... but we don't know where it starts. We would need to know either how far from the side of the bridge it's start point is, or at least it's trajectory...

OpenStudy (anonymous):

it is probably to late but this is what I got... Let X denote the vertical distance that it spans. Then, since it spans 1 more foot horizontally, it spans x+2 feet horizontally. x^2+(x+2)^2=10^2 =>x^2+(x^2+4x+4)=100 by expanding =>2x^2+4x+4=100 =>2x^2+4x-96=0, by setting the right side equal to zero. =>2(x+8)(x-6)=0, by factoring =>x=-8 and x=6 by the zero-product property Since length cannot be negative, x=6. Therefore, the brace spans 6 ft vertically and 6+2=8 ft horizontally I got some help with this problem, so I believe it is correct... Hope this helps! good luck